OpenStudy (anonymous):

A student visiting the Sears Tower Skydeck is 1353 feet above the ground. Find the distance the student can see to the horizon. Use the formula to approximate the distance d in miles to the horizon when h is the height of the viewer’s eyes above the ground in feet. Round to the nearest mile. A. 45 miles B. 601 miles C. 36 miles D. 1010 miles

OpenStudy (ashleyisakitty):

Did it include what d equals to? d=?

OpenStudy (anonymous):

\[d=\sqrt{1.5h}\]

OpenStudy (ashleyisakitty):

d = (1.5*1353)^1/2 d = 2029.5^1/2 d = 45.04

OpenStudy (anonymous):

THANK YOU!!!

OpenStudy (ashleyisakitty):

no problem :)